OpenStudy (habbababba1):

please help will fan and medal...Express each quotient in the simplest form.

OpenStudy (habbababba1):

im not to good at math what exactly does it mean that they r functions of x all the answer choices are y and z choices

OpenStudy (habbababba1):

its fine i just got confused

OpenStudy (mathstudent55):

Take the given expression and do the actual division. To divide a fraction by a fraction, multiply the first fraction by the reciprocal of the second fraction. \(\large \dfrac{10xy^2}{3z} \div \dfrac{5xy}{6z^3} =\) Multiply the first fraction by the reciprocal of the second fraction. \(\large =\dfrac{10xy^2}{3z} \times \dfrac{6z^3} {5xy} \) To multiply fractions, multiply the numerators together, and multiply the denominators together. \(\large =\dfrac{10xy^2 \times6z^3 }{3z \times 5xy} \) \(\large =\dfrac{60xy^2z^3 }{15xyz} \) The last step is to reduce, which I will leave for you to do.